Here is a table of zero coupon bond prices:

Maturity: 1 year 2 year 3 year

Price: 0.99010 0.97066 0.0.94929

a) What are the corresponding spot rates (1yr, 2yr, 3yr)?

b) What is the 1 year rate deferred 1 year F[1,2]?

c) What is the 1 year rate deferred 2 years F[2,3].

d) What is the 2 year rate deferred 1 year F[1,3]?
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

B C Maturity in Year -MO00 Bond Price 0.99010 0.97066 0.94929 D (a) Spot rate 1.00% 1.50% 1.75% 3 F[1,2] F[2,3] 7 b) 8 c) 9 d

Cell reference -

2 A B Maturity in Year Bond Price (a) Spot rate =(1/3)^(1/B3)-1 =(1/C4)^(1/84)-1 =(1/C5)^(1/B5)-1 0.9901 0.97066 0.94929 4 7

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.